OpenStudy (anonymous):

1. Cost of Tuition In 2000, the average cost of tuition and fees at private four-year colleges was $16,200, and in 2005 it was $20,100. Sketch a line that passes through the points (2000, 16200) and (2005, 20100). (Source: The College Board.) 2. Rate of Change in Tuition Calculate the slope of the line in your graph. Interpret this slope as a rate of change. 3. Modeling Tuition Find the slope–intercept form of the line in your sketch. What is the y-intercept and does it have meaning in this situation? 4. Predicting Tuition Use your equation to estimate tuition and fees in 2001 and compare it

OpenStudy (blurbendy):

for the first one, label your x axis as years and your y axis as tuition amount. then plot the points

OpenStudy (blurbendy):

for the second one, slope = (y2 - y1) / (x2 - x1)

OpenStudy (blurbendy):

tell me what you get

OpenStudy (anonymous):

the rate of the slope is 3/5

OpenStudy (blurbendy):

where did you get the 3 from?

OpenStudy (anonymous):

the cordinates are (0,1) (5,4) right?

OpenStudy (blurbendy):

your coordinates should be: (2000, 16200) and (2005, 20100)

OpenStudy (anonymous):

i mean the slope of the line sorry

OpenStudy (anonymous):

is 3/5

OpenStudy (blurbendy):

the slope and the rate of change are the same thing. to get the slope, you would calculate: (21,600 - 16200) / (2005 - 2000)

OpenStudy (anonymous):

im confused with how to solve part four

OpenStudy (blurbendy):

solve part 2 first

OpenStudy (blurbendy):

it will help you solve part 4

OpenStudy (blurbendy):

did you get an answer?

OpenStudy (anonymous):

5/3900

OpenStudy (blurbendy):

let's break it down. what is (21,600 - 16200)

OpenStudy (anonymous):

3900

OpenStudy (blurbendy):

no, it's 5,400

OpenStudy (anonymous):

oops sorry pushin wrong button

OpenStudy (blurbendy):

so we have 5,400 / (2005 - 2000) which is: 5,400 / 5 = 1080

OpenStudy (blurbendy):

so our slope is 1080, understand?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

okay, for 3 we need it in slope intercept form: y - y1 = 1080(x - x1) let's use (2000, 16200) and put them in for y1 and x1 respectively y - 16200 = 1080(x - 2000) y - 16200 = 1080x -2160000 y = 1080x - 2143800

OpenStudy (anonymous):

im sorry im just having so much trouble with this whole problem. i get this, but step four, im just lost

OpenStudy (blurbendy):

okay, we're simply just going to put 2001 in for x in our equation. y = 1080(2001) - 2143800 y = 17280 that was the cost of tuition in 2001 which was more $1080 more than in 2000 and 2820 less than 2005

OpenStudy (anonymous):

Thank you so much, i am trying to re read through it all to really understand it

OpenStudy (blurbendy):

anytime! let me know if you have any questions

OpenStudy (anonymous):

thank you again

OpenStudy (blurbendy):

my pleasure